# On a class of interdiction problems with partition matroids: complexity and polynomial-time algorithms

### v4

@@ 2.1 Integer programming formulation
$L_{k}\subset N$ and $l_{k}\in\mathbb{Z}_{+}$, $k\in K_{l}$ leader’s partition of $N$ and respective budgets of the leader
$F_{k}\subset N$ and $f_{k}\in\mathbb{Z}_{+}$, $k\in K_{f}$ follower’s partition of $N$ and respective budgets of the follower
@@ 2.2 Proof of $NP$-hardness
~$\widetilde{\mathcal{I}}_{X}=\Big{\{}\mathbf{x}\in\{0,1\}^{|V|\times|E|}:\sum_{v\in V}x_{v,e}\leq 1\quad\forall e\in E\Big{\}}\,\mbox{ and }$ (3a)
~$\widetilde{\mathcal{I}}_{Y}(\mathbf{x})=\Big{\{}\mathbf{y}\in\{0,1\}^{|V|\times|E|}:\;\mathbf{y}\leq 1-\mathbf{x},\;\sum_{e\in E}y_{v,e}\leq 1\quad\forall v\in V\Big{\}},$ (3b)
@@
$x_{v,e}=\begin{cases}1,\mbox{ if }\widetilde{\beta}_{v,e}=1\mbox{ and }v\in S,\\ 0,\mbox{ otherwise.}\\ \end{cases}$
@@ 3.1 Duality-based algorithm
In particular, the integrality constraints $\mathbf{y}\in\{0,1\}^{n}$ can be relaxed, since any matroid polyhedron, i.e., a convex hull of incidence vectors of its independent sets, is integral; see, e.g., [17, 41]. Furthermore, if superscript “$(k)$” for $k\in K_{f}$ refers to the subset of indices $F_{k}$, then (6) can be expressed in a block-diagonal form as:
@@
Then, we observe that objective function (9a) is bilinear in the dual variables $\gamma_{i}$ and the upper-level decision variables $x_{i}$, $i\in F_{k}$. In order to linearize the resulting single-level reformulation, we divide the interval $[0,\beta_{n}]$ induced by constraints $\alpha_{k}\geq 0$ and $\alpha_{k}\leq\beta_{n}$ into $n$ disjoint intervals
$[\beta_{j},\beta_{j+1}],\;j\in\{0,\ldots,n-1\}.$
@@
Here, we exploit the fact that the overall minimum on the left-hand side of (11) is taken with respect to $2n$ terms, $n-1$ of which are duplicated. Eventually, the dual reformulation of [PMI] is obtained by combining the minimum over $\mathbf{x}\in\mathcal{I}_{X}$ and the right-hand side of (11). ∎
@@ 3.2 Dynamic programming-based algorithm
and the superscript “$(k)$” for $k\in K_{f}$ refers to the subset of indices $F_{k}$. Put differently, [PMI] can be divided into $k_{f}$ distinct stages, where at each stage $k\in K_{f}$ we need to identify an optimal blocking decision $\mathbf{x}^{*\text{\tiny({k})}}$ in the subset $F_{k}$ controlled by the follower.
@@ 3 $k\longleftarrow k_{f}$
6       $\mathbf{x}^{*\text{\tiny({k})}}(\mathbf{r}_{k})\in\operatorname{argmin}_{\,\mathbf{x}^{\text{\tiny({k})}}\in A^{\text{\tiny({k})}}(\mathbf{r}_{k})}\psi_{k}(\mathbf{x}^{\text{\tiny({k})}})$
@@
12             $\boldsymbol{\theta}_{k}(\mathbf{r}_{k},\mathbf{x}^{\text{\tiny({k})}})\longleftarrow\left(r_{k,k^{\prime}}-\sum_{i\in F_{k}\cap L_{k^{\prime}}}x_{i}\right)_{k^{\prime}\in K_{l}}$
@@
14             $\mathbf{x}^{*\text{\tiny({k})}}(\mathbf{r}_{k})\in\arg\min_{\,\mathbf{x}^{\text{\tiny({k})}}}\Big{\{}\psi_{k}(\mathbf{x}^{\text{\tiny({k})}})+V_{k+1}\big{(}\boldsymbol{\theta}_{k}(\mathbf{r}_{k},\mathbf{x}^{\text{\tiny({k})}})\big{)}:\mbox{ }\mathbf{x}^{\text{\tiny({k})}}\in A^{\text{\tiny({k})}}(\mathbf{r}_{k})\Big{\}}$
15
@@
20 $\mathbf{x}^{\text{\tiny(1)}}\longleftarrow\mathbf{x}^{\text{\tiny(1)}}(\mathbf{r}^{*}_{1})$
@@
23       $\mathbf{x}^{*\text{\tiny({k})}}\longleftarrow\mathbf{x}^{*\text{\tiny({k})}}(\mathbf{r}^{*}_{k})$
@@
A pseudocode of the proposed DP-based algorithm is illustrated in Algorithm 1. Initially, at each stage $k\in\{k_{f},\ldots,1\}$, we compute a value function $V_{k}(\mathbf{r}_{k})$, which represents the minimum total loss the leader can achieve from a given state $\mathbf{r}_{k}\in R_{k}$ by following an optimal interdiction strategy; see lines 5 and 11 of Algorithm 1. In particular, in line 10, we express a new state, $\mathbf{r}_{k+1}$, as a function of $\mathbf{r}_{k}$ and $\mathbf{x}^{\text{\tiny({k})}}$ by using equation (16). In lines 6 and 12, for each state $\mathbf{r}_{k}\in R_{k}$, we also identify an optimal leader’s decision, $\mathbf{x}^{*\text{\tiny({k})}}(\mathbf{r}_{k})$. It is noteworthy that, according to lines 6 and 12, Algorithm 1 identifies some optimal solution of [PMI] even when multiple optimal solutions exist. Finally, an optimal blocking decision, $\mathbf{x}^{*}$, is restored in lines 15–20 of Algorithm 1.
Overall, Algorithm 1 is based on a standard backward dynamic programming approach, operating over a finite state space and a finite action space, as defined by equations (14) and (15), respectively. Consequently, its correctness follows from Bellman’s optimality principle [4]. The following result establishes the computational complexity of Algorithm 1.
@@
for each $\mathbf{r}_{k+1}\in R_{k+1}$ is precomputed at stage $k+1$. Finally, from lines 1414141420202020 of Algorithm 1 we observe that an optimal solution $\mathbf{x}^{*}$ can be restored in $O(n)$ operations.
@@ 3 $\mathbf{x}\longleftarrow\mathbf{0}$
4 $N(\mathbf{x})$longleftarrow $\{i\in N:x_{i}=0\,\mbox{ and }\,\mathbf{x}+\mathbf{e}_{i}\in\mathcal{I}_{X}\}$
@@ 7       $\mathbf{x}\longleftarrow\mathbf{x}+\mathbf{e}_{i^{*}}$
$N_{+}(\mathbf{x}):=\Big{\{}i\in N(\mathbf{x}):\,i\geq j(\mathbf{x},f)\Big{\}}\,\mbox{ and }\,N_{-}(\mathbf{x}):=\Big{\{}i\in N(\mathbf{x}):\,i<j(\mathbf{x},f)\Big{\}}.$
@@ 4 Conclusion
and the superscript “$(k)$” for $k\in K_{f}$ refers to the subset of indices $F_{k}$. Clearly, the increase of $\mathbf{x}$ can only reduce the follower’s feasible sets in (26) and, therefore, it cannot increase the leader’s objective function value, $\psi(\mathbf{x})$. Hence, $\psi(\mathbf{x})$ is monotone non-increasing.

### v5

@@ 2.1 Integer programming formulation
$L_{k}\subseteq N$ and $l_{k}\in\mathbb{Z}_{+}$, $k\in K_{l}$ leader’s partition of $N$ and respective budgets of the leader
$F_{k}\subseteq N$ and $f_{k}\in\mathbb{Z}_{+}$, $k\in K_{f}$ follower’s partition of $N$ and respective budgets of the follower
@@ 2.2 Proof of $NP$-hardness
$\widetilde{\mathcal{I}}_{X}=\Big{\{}\mathbf{x}\in\{0,1\}^{|V|\times|E|}:\sum_{v\in V}x_{v,e}\leq 1\quad\forall e\in E\Big{\}}\,\mbox{ and }$ (3a)
$\widetilde{\mathcal{I}}_{Y}(\mathbf{x})=\Big{\{}\mathbf{y}\in\{0,1\}^{|V|\times|E|}:\;\mathbf{y}\leq 1-\mathbf{x},\;\sum_{e\in E}y_{v,e}\leq 1\quad\forall v\in V\Big{\}},$ (3b)
@@
[…]
@@ 3.1 Duality-based algorithm
In particular, the integrality constraints $\mathbf{y}\in\{0,1\}^{n}$ can be relaxed, since any matroid polyhedron, i.e., a convex hull of incidence vectors of its independent sets, is integral; see, e.g., [17, 41]. Furthermore, if superscript “(k)(k)” for $k\in K_{f}$ refers to the subset of indices $F_{k}$, then (6) can be expressed in a block-diagonal form as:
@@
Then, we observe that objective function (9a) is bilinear in the dual variables $\gamma_{i}$ and the upper-level decision variables $x_{i}$, $i\in F_{k}$. In order to linearize the resulting single-level reformulation, we divide the interval $0,\beta_{n}]$ induced by constraints $\alpha_{k}\geq 0$ and $\alpha_{k}\leq\beta_{n}$ into $n$ disjoint intervals
$\beta_{j},\beta_{j+1}],\;j\in\{0,\ldots,n-1\}.$
@@
Here, we exploit the fact that the overall minimum on the left-hand side of (11) is taken with respect to $2n$ terms, n1n-1 of which are duplicated. Eventually, the dual reformulation of [PMI] is obtained by combining the minimum over $\mathbf{x}\in\mathcal{I}_{X}$ and the right-hand side of (11). ∎
@@ 3.2 Dynamic programming-based algorithm
is the follower’s profit associated with the subset $F_{k}$ under a blocking decision $\mathbf{x}^{\text{\tiny({k})}}$, and the superscript “(k)(k)” for $k\in K_{f}$ refers to the subset of indices $F_{k}$. Put differently, [PMI] can be divided into $k_{f}$ distinct stages, where at each stage $k\in K_{f}$ we need to identify an optimal blocking decision $\mathbf{x}^{*\text{\tiny({k})}}$ in the subset $F_{k}$.
@@ 3 $k\longleftarrow k_{f}$
6  $\tilde{\mathbf{x}}^{*\text{\tiny({k})}}(\mathbf{r}_{k})\in\operatorname{argmin}_{\,\mathbf{x}^{\text{\tiny({k})}}\in A^{\text{\tiny({k})}}(\mathbf{r}_{k})}\psi_{k}(\mathbf{x}^{\text{\tiny({k})}})$
@@
13     $\tilde{\mathbf{x}}^{*\text{\tiny({k})}}(\mathbf{r}_{k})\in\arg\min_{\,\mathbf{x}^{\text{\tiny({k})}}}\Big{\{}\psi_{k}(\mathbf{x}^{\text{\tiny({k})}})+V_{k+1}\big{(}\boldsymbol{\theta}_{k}(\mathbf{r}_{k},\mathbf{x}^{\text{\tiny({k})}})\big{)}:\mbox{ }\mathbf{x}^{\text{\tiny({k})}}\in A^{\text{\tiny({k})}}(\mathbf{r}_{k})\Big{\}}$
14
@@
19 $\mathbf{x}^{*\text{\tiny(1)}}\longleftarrow\tilde{\mathbf{x}}^{*\text{\tiny(1)}}(\mathbf{r}^{*}_{1})$
@@
22  $\mathbf{x}^{*\text{\tiny({k})}}\longleftarrow\tilde{\mathbf{x}}^{*\text{\tiny({k})}}(\mathbf{r}^{*}_{k})$
@@
A pseudocode of the proposed DP-based algorithm is illustrated in Algorithm 1. Initially, at each stage $k\in\{k_{f},\ldots,1\}$, we compute a value function $V_{k}(\mathbf{r}_{k})$, which represents the minimum total loss the leader can achieve from a given state $\mathbf{r}_{k}\in R_{k}$ by following an optimal interdiction strategy; see lines 5 and 10 of Algorithm 1. In lines 6 and 11, for each state $\mathbf{r}_{k}\in R_{k}$, we also identify an optimal leader’s decision, $\tilde{\mathbf{x}}^{*\text{\tiny({k})}}(\mathbf{r}_{k})$, within the subset $F_{k}$. It is noteworthy that, according to lines 6 and  11, Algorithm 1 identifies some optimal solution of [PMI] even when multiple optimal solutions exist. Finally, an optimal blocking decision, $\mathbf{x}^{*}$, is restored in lines 14–19 of Algorithm 1.
Overall, Algorithm 1 is based on a standard backward dynamic programming approach, operating over a finite state space and a finite action space, as defined by equations (14) and (15), respectively. In particular, the value functions in lines 5 and 10 of Algorithm 1 are computed using the standard recursive formula of dynamic programming. Consequently, correctness of Algorithm 1 follows from Bellman’s optimality principle [4]. The following result establishes the computational complexity of Algorithm 1.
@@
for each $\mathbf{r}_{k+1}\in R_{k+1}$ is precomputed at stage $k+1$. Finally, from lines 14142020 of Algorithm 1 we observe that an optimal solution $\mathbf{x}^{*}$ can be restored in $O(n)$ operations.
@@ 7  $\mathbf{x}\longleftarrow\mathbf{x}+\mathbf{e}_{i^{*}}$
[…]
@@
Notably, the proof of Theorem 4 shows that [PMI] with $k_{f}=1$ can be solved by leveraging the maximum weight independent set problem in the leader’s matroid $X$. However, Algorithm 2 leverages the structure of [PMI] with $k_{f}=1$ and, depending on the tie-breaking rule in line 6, can produce other optimal solutions that differ from the maximum weight independent set in $X$.
@@ 4 Conclusion
and the superscript “(k)(k)” for $k\in K_{f}$ refers to the subset of indices $F_{k}$. Clearly, the increase of $\mathbf{x}$ can only reduce the follower’s feasible sets in (26) and, therefore, it cannot increase the leader’s objective function value, $\psi(\mathbf{x})$. Hence, $\psi(\mathbf{x})$ is monotone non-increasing.